# Vintern-1B: An Efficient Multimodal Large Language Model for Vietnamese

### v1

@@ Abstract
In this report, we introduce Vintern-1B (Vietnamese-InternVL2-1B), a reliable 1-billion-parameters multimodal large language model (MLLM) for Vietnamese language tasks. By integrating the Qwen2-0.5B-Instruct [1] language model with the InternViT-300M-448px [2] visual model, Vintern-1B is optimized for a range of applications, including optical character recognition (OCR), document extraction, and general question-answering in Vietnamese context. The model is fine-tuned on an extensive dataset of over 3 million image-question-answer pairs, achieving robust performance and reliable results across multiple Vietnamese language benchmarks like OpenViVQA [3]. Vintern-1B is small enough to fit into various on-device applications easily. Additionally, we have open-sourced several Vietnamese vision question answering (VQA) datasets for text and diagrams, created with Gemini 1.5 Flash [4]. Our models are available at: https://huggingface.co/5CD-AI/Vintern-1B-v2.
@@ 1 Introduction
To address these limitations, we introduce Vintern-1B, an efficient 1-billion-parameter multimodal model specifically designed for Vietnamese language tasks. Vintern-1B based on InternVL2-1B [2], which integrates the Qwen2-0.5B-Instruct [1] language model with the InternViT-300M-448px [2] visual model, optimized for tasks such as OCR, document extraction, and general VQA. Additionally, we have created new OCR and document visual question answering (DocVQA) datasets for the Vietnamese language, which have been used to fine-tune Vintern-1B. We provide a detailed description of Vintern-1B’s architecture, data creation, training procedures, and how to benchmark MLLMs using MLLM-as-a-Judge.
@@ 2 Related work
The exceptional performance of GPT-4V [8] extends GPT-4’s capabilities by incorporating visual inputs, allowing it to handle both text and image content, marking a significant advancement in the domain of Multimodal Large Language Models (MLLMs). Following this, Gemini Pro Vision [14] further enhances MLLMs by enabling the processing of text, images, and audio, while supporting up to 1 million tokens, significantly boosting performance.
Over the past year, many open-source MLLMs have gained prominence. Recent research has focused on developing MLLMs that achieve unified understanding and reasoning across different modalities, building on the success of LLMs. Models such as the LLaVA series, MiniGPT-4, Qwen-VL, Florence-2, and the InternVL series have demonstrated impressive results by aligning visual input with LLMs through visual instruction tuning.
In the context of Vietnamese language, research efforts such as V-Vistral [11] and LaVy [13] have made significant contributions to the progress of MLLMs. However, these models still exhibit notable gaps in understanding documents, charts, infographics, and recognizing scene texts within the Vietnamese context, compared to other open-source models in English or Chinese. This disparity is largely due to the lack of high-quality Vietnamese datasets for MLLMs.
In this report, we address this challenge by introducing Vintern-1B, an effective MLLM model for Vietnamese, fine-tuned on our large-scale, high-quality datasets. Additionally, we benchmark the model using MLLMs-as-judge, specifically designed to evaluate MLLMs’ understanding of Vietnamese visual language tasks.
@@ 3.1 Overall Architecture
Vintern-1B model are built with InternVL-1.5 [2] architecture akin to widely-used open-source MLLMs, specifically the “ViT-MLP-LLM” configuration referenced in various existing studies, see detailed in Figure 1. The architecture using three main components:
  • Vision Encoder: InternViT‑300M‑448px [2] a distilled small vision foundation model, is used as the vision encoder.
  • MLP Projector: A two-layer Multi-Layer Perceptron (MLP) projector is employed to align the output representations from the visual and language modalities. This projector ensures that the visual and textual information is transformed into a common space.
  • Large Language Model: We utilize the pre-trained Qwen2‑0.5B‑Instruct [1], which is a language model responsible for generating textual information, take the aligned representations from the MLP projector
During training, the model uses module dynamic resolution strategy, dividing images into tiles of 448×448 pixels in sizes ranging from 1 to 12, based on the aspect ratio and resolution of the input images. During testing, this can be zero-shot scaled up to 40 tiles (i.e., 4K resolution). To enhance scalability for high resolution, a pixel shuffle operation is employed to reduce the number of visual tokens to one-quarter of the original. Therefore, in the model, a 448×448 image is represented by 256 visual tokens.
Figure 1: Overall Architecture. Vintern-1B adopts the ViT-MLP-LLM architecture similar to popular MLLMs ([2] [15] [16] [6]), combining a pre-trained InternViT‑300M‑448px [2] with Qwen2‑0.5B‑Instruct [1] through a MLP projector. The model also employ a simple pixel shuffle to reduce the number of visual tokens to one-quarter. Module Dynamic High Resolution dynamically match an optimal aspect ratio from pre-defined ratios, dividing the image into tiles of 448×448 pixels and creating a thumbnail for global context. This method minimizes aspect ratio distortion and accommodates varying resolutions during training.
@@ 3.2.3 General Question Answering (QA)
  • Viet-OpenViVQA-gemini-VQA based on OpenViVQA [3]: This dataset focuses on open-domain question answering, providing the model with exposure to various types of queries, from straightforward factual questions to more complex, context-dependent ones.
  • Viet-Localization-VQA includes quintessentially Vietnamese images such as scenic landscapes, historical sites, culinary specialties, festivals, cultural aspects from various regions, familiar rural scenes, and everyday life in urban areas, among others.
@@ 4.1 Implementation Details
Vintern-1B was developed by integrating the InternViT‑300M‑448px [2] vision encoder with the Qwen2‑0.5B‑Instruct [1] language model, using a dynamic high-resolution strategy. In this approach, images are segmented into 448×448 pixel tiles, with the number of tiles ranging up to 12 based on the image’s aspect ratio and resolution during training. In testing phases, the model could handle up to 40 tiles, equivalent to 4K resolution, demonstrating its adaptability to highresolution inputs in a zero-shot manner.
@@
The initial version of Vintern-1B-v1 was fine-tuned from InternVL2-1B using the Viet-Doc-VQA, Viet-Doc-VQA-II OCR and Viet-OCR-VQA datasets for 1 epoch. In the next phase, Vintern-1B-v2, was further fine-tuned by continuing from v1, using the same Viet-Doc-VQA, Viet-Doc-VQA-II OCR, Viet-OCR-VQA and datasets along with all remaining datasets in 3.2 for an additional 1 epoch. Both versions were trained on 4 GPUs (Nvidia RTX-3090) with a global batch size of 128, a learning rate of 4e-5, and a context length of 4096. We adopted the same response formatting prompts as LLaVA 1.5 [16].

### v2

@@ Abstract
In this report, we introduce Vintern-1B (Vietnamese-InternVL2-1B), a reliable 1-billion-parameters multimodal large language model (MLLM) for Vietnamese language tasks. By integrating the Qwen2-0.5B-Instruct [1] language model with the InternViT-300M-448px [2] visual model, Vintern-1B is optimized for a range of applications, including optical character recognition, document information extraction, and general visual question-answering in Vietnamese context. The model is fine-tuned on an extensive dataset of over 3 million image-question-answer pairs, achieving robust performance and reliable results across multiple Vietnamese language benchmarks like OpenViVQA [3]. Vintern-1B is small enough to fit into various on-device applications easily. Additionally, we have open-sourced several Vietnamese vision question answering (VQA) datasets for text, documents, diagrams, and more created with Gemini 1.5 Flash [4]. Our models are available at: https://huggingface.co/5CD-AI/Vintern-1B-v2.
@@ 1 Introduction
To address these limitations, we carefully created datasets with a focus on text and document processing, as well as additional datasets featuring Vietnamese localization images. We then fine-tuned the InternVL2-1B model, which has achieved impressive results in text processing with our datasets. We provide a detailed description of Vintern-1B’s architecture, data creation, training procedures, and how to benchmark our MLLMs.
@@ 2 Related work
2.1 Multimodal Large Language Model
Closed-source MLLMs, such as GPT-4o [14] and Gemini Pro Vision [15], have significantly advanced the integration of visual inputs, enabling sophisticated handling of text, images, and audio. Concurrently, the past year has witnessed a surge in open-source MLLMs, where researchers have been striving to achieve unified reasoning across modalities by leveraging the advancements in LLMs. A notable contribution in this space is the InternVL series [5] [2], which has showcased impressive results by aligning visual inputs with LLMs through visual instruction tuning.
InternVL2-1B [2], for instance, employs the InternViT-300M-448px, a distilled vision foundation model with 300M parameters derived from the robust vision encoder InternViT-6B-448px, which possesses continuous learning capabilities. This enhancement significantly boosts its visual understanding, enabling seamless transfer and reuse across different LLMs. The language model Qwen2-0.5B-Instruct [1], which has been integrated into this framework, demonstrates competitive multilingual capabilities, proficient in approximately 30 languages, particularly excelling in Vietnamese, promising strong support for MLLMs designed for the Vietnamese language.
2.2 Vietnamese MLLMs
In the Vietnamese language context, recent models like V-Vistral [11] and LaVy [13] have made notable progress in advancing MLLMs. However, these models still struggle with understanding documents, charts, infographics, and recognizing scene texts specific to Vietnamese. While large Vietnamese MLLM datasets like Vistra [12], which was translated from English to Vietnamese using Gemini Pro, have enhanced the naturalness of question-answering, the overall model performance remains suboptimal. This disparity is mainly due to the lack of high-quality, purely Vietnamese image content. The situation is further complicated by the absence of diverse datasets that support various tasks such as OCR and information extraction, which are essential for improving the capabilities of MLLMs in Vietnamese.
@@ 3.1 Overall Architecture
The Vintern-1B model is constructed using the InternVL-1.5 1 architecture, which is similar to the ’ViT-MLP-LLM’ configuration commonly found in many open-source MLLMs, as discussed in several studies. For more details, refer to Figure 1.
  • Vision Encoder: InternViT‑300M‑448px [2] a distilled small vision foundation model, is used as the vision encoder, which transform the image to visual features.
  • MLP Projector: A two-layer Multi-Layer Perceptron (MLP) projector is utilized to map the output representations from both the visual and language modalities into a unified space, facilitating the integration of visual and textual information.
  • Large Language Model: We utilize the pre-trained Qwen2‑0.5B‑Instruct [1] serves as our language model, which generates text by leveraging the aligned representations provided by the MLP projector.
Our training processs follows the InternVL 1.5 [2], where images are segmented into $448\times 448$ pixel tiles. We use pixel shuffle to reduce token count, enabling efficient high-resolution scaling, each $448\times 448$ image is encoded as 256 visual tokens within the model.
Figure 1: Overall Architecture. Vintern-1B is built upon the ViT-MLP-LLM framework, following the structure of well-known MLLMs ([2] [16] [17] [6]). It inherits from InternVL 1.5 [2], integrating a pre-trained InternViT-300M-448px [2] with Qwen2-0.5B-Instruct [1] via an MLP projector. The input image is processed by the Dynamic High Resolution module, which splits it into smaller 448x448 pixel images along with a thumbnail. These images are then passed through InternViT-300M-448px to extract visual features. A Pixel Shuffle step is also applied before feeding the data into the MLP projector to align it with the embeddings of the large language model Qwen2-0.5B-Instruct which takes the aligned visual tokens and the related question as inputs, and generates the corresponding answer.
@@ 3.2.3 General Question Answering (QA)
  • Viet-OpenViVQA-gemini-VQA: based on OpenViVQA [3]: This dataset focuses on open-domain question answering, providing the model with exposure to various types of queries, from straightforward factual questions to more complex, context-dependent ones.
  • Viet-Localization-VQA: includes quintessentially Vietnamese images such as scenic landscapes, historical sites, culinary specialties, festivals, cultural aspects from various regions, familiar rural scenes, and everyday life in urban areas, among others.
@@ 4.1 Implementation Details
Vintern-1B employed a dynamic high-resolution approach, where images were divided into 448×448 pixel tiles. The number of tiles could vary, reaching up to 12 based on the aspect ratio and resolution during training. In the testing phase, the model could handle as many as 12 tiles.
@@
The initial version of Vintern-1B-v1 was fine-tuned from InternVL2-1B using the Viet-Doc-VQA, Viet-Doc-VQA-II OCR and Viet-OCR-VQA datasets for 1 epoch. In the next phase, Vintern-1B-v2, was further fine-tuned by continuing from v1, using the same Viet-Doc-VQA, Viet-Doc-VQA-II OCR, and Viet-OCR-VQA datasets along with all remaining datasets in 3.2 for an additional 1 epoch. Both versions were trained on 4 GPUs (Nvidia RTX-3090) with a global batch size of 128, a learning rate of 4e-5, and a context length of 4096. Our approach utilizes a context length of 4096, and we follow the response formatting prompts outlined in LLaVA 1.5 [17].